Jeramie has a jar of nickels and dimes. There are 200 coins worth $14.00. How many of each type of coin are in the jar?
Answer by ptaylor(2198)

Letx=number of nickels
Then 200-x=number of dimes

Now lets deal in pennies to keep down confusion.
We are told that:
5x+10(200-x)=1400 all in pennies; divide each term by 5
x+2(200-x)=280 clear parens
x+400-2x=280 subtract 400 from both sides
-x=-120
x=120----------------------number of nickels
200-x=200-120=80-------------number of dimes
ck
120(5)+80(10)=1400
600+800=1400
1400=1400

Hope this helps----ptaylor